OpenStudy (anonymous):

The electric potential in a circuit is given by V(t) = 120−120e^kt, where k is an unknown constant. Assume that t is measured in seconds and V is measured in volts. If V(0.5) = 66.1V, when will V(t) = 100V? Be accurate to two decimal places

OpenStudy (anonymous):

You have that: \(120(1 - e^{kt}) = V(t)\) Now our unknown constant can be solved for: \(120(1 - e^{k/2}) = 66.1 \\ 1 - e^{k/2} = \frac{66.1}{120} \\ e^{k/2} = -\frac{66.1}{120} + 1 \\ \frac k 2 = \ln (-\frac{66.1}{120} + 1) k = 2\ln (-\frac{66.1}{120} + 1)\)

OpenStudy (anonymous):

E.g., k = 2log(-55.1/120 + 1), which is about -1.22.

OpenStudy (anonymous):

Substitute k into your original equation and V(t) = 100, and solve for t.

OpenStudy (anonymous):

oh okay, I don't know why I didn't see that before thank you for the help. It makes a lot of sense now!